5power of3÷(10+15)+9power2+33
Answer by stanbon(75887)

5power of3÷(10+15)+9power2+33
----
= 5^3 119/ [25] + 9^2 + 33
----
= 125/25 + 81 + 33
----
= 5 + 81 + 33
----
= 119
===========
Cheers,
Stan H.
===========